# Reinforcing Infrastructure Networks with Multicriteria Portfolio Decision Analysis: An Application to Railway Stations in Finland

### v1

@@ 4.2 Preferences Regarding Connections
We model two situations regarding the preferences. In the first, the DM does not state preferences about the connections. This situation is represented by the set $\mathcal{S}^{0}=\big{\{}w\in\mathbb{R}^{3}\mid w_{j}\geq 0\ \forall j,\sum_{j=1}^{3}w_{j}=1\big{\}}$.
In the second, the DM establishes a ranking of the connections based on their yearly traffic volume: (A,B) at 500 trains/year, (A,C) at 500 trains/year, and (B,C) at 100 trains/year. This is represented by the set $\mathcal{S}^{1}=\big{\{}w\in\mathbb{R}^{3}\mid\;w_{j}\geq 0\ \forall j,\;\sum_{j=1}^{3}w_{j}=1,w_{2}\geq 5w_{1},\;w_{3}\geq 5w_{1}\big{\}}$. For this last situation, we study what happens if the DM aims to keep the performance of the individual connections above 95% and 96%. We denote the cases with performance requirements by $\mathcal{S}^{1}\wedge\alpha=0.95$ and $\mathcal{S}^{1}\wedge\alpha=0.96$ respectively.
@@ 4.3.1 Cost-efficient Portfolios
We consider a maximum of 20 reinforcement actions (i.e., it is possible to reinforce up to 20 switches). The reinforcement cost is the same for each switch. Applying Algorithm 1, we identify the cost-efficient portfolios in Table 1 for different information sets and minimum reliability requirements. As a reference, the number of feasible portfolios is presented for different numbers of reinforced switches.
Table 1: Number of cost-efficient portfolios by information set and reliability requirements.
Reinforced switches 1 5 10 15 20
@@
When preferences about the traffic volume and the minimum reliability requirements are provided, the number of cost-efficient portfolios becomes much smaller. These results highlight the benefits of eliciting preferences regarding the connectivity between borderline nodes in narrowing down the reinforcement alternatives and easing the selection process.
@@ 4.3.2 Reliability of the Connections
The reliability of the connections as a result of implementing the cost-efficient portfolios for the set $S^{0}$ is in Figure 3. Each point corresponds to a different portfolio for a given connection. The marginal contribution to improve the reliability of a connection (i.e., the increase in the probability of the connection being available when reinforcing an additional switch) decreases with the number of reinforced switches. This is because the resources are allocated first to the more beneficial actions, and we consider actions of equal cost. Otherwise, it is possible to have an increasing marginal contribution to reducing failures because more resources allow not just the implementation of more actions but also the inclusion of actions that are not affordable with fewer resources.
@@ 4.3.3 Minimum Performance for a Single Connection
The reliability of the connections as a result of implementing each cost-efficient portfolio for up to five reinforced switches is in Figure 4 for the non-informative and the incomplete information set. Given that the traffic volume between nodes B and C is smaller than in the other two connections, when including the preferences based on the traffic volume (incomplete set), the connection (B, C) performs worse as a result of implementing the cost-efficient portfolios compared to the information set with no stated preferences.
The non-informative set contains the weights that give importance only to a single connection: (1,0,0), (0,1,0), and (0,0,1). Therefore, the maximum reliability achievable at each connection can be obtained by implementing a cost-efficient portfolio in this set (see Lemma 1). These maximum reliability values at each connection can support resource allocation decisions by quantifying the potential improvements in reliability for a given budget. For example, the maximum reliability of the connection (B, C) achieved by reinforcing up to five switches is around $0.96$.
Moreover, if there is a minimum performance requirement such that the reliability of a connection cannot be less than $0.95$, not all the cost-efficient portfolios in the incomplete information set $S^{1}$ satisfy it for the path (B, C). Suppose the requirement is higher, such as 0.96. In that case, none of the cost-efficient portfolios in the incomplete set fulfills it, even though some feasible portfolios that are not cost-efficient $S^{1}$ can comply with the requirement.
We apply Algorithm 2, enforcing a minimum performance of 0.95 and 0.96. In addition to computing the cost-efficient portfolios (two last rows of Table 1), this analysis also provides information about the minimum number of switches that must be reinforced to fulfill minimum reliability requirements. For instance, at least three switches must be reinforced to comply with the 0.95 requirements and five for the 0.96 requirements.
Figure 4: Reliability of connections for cost-efficient portfolios containing up to five reinforced switches. Values are presented for an incomplete information set ($S^{1}$) and the non-informative set ($S^{0}$).
4.3.4 Selecting Switches to Reinforce
@@
$CI(q_{k},c)=\dfrac{\lvert\{q\in\mathcal{Q}_{ND}(c)\mid q_{k}\in q\}\rvert}{\lvert\mathcal{Q}_{ND}(c)\rvert}$ (6)
If the budget has been determined, a reinforcement action with a core index of 1 can be recommended because it is in every non-dominated portfolio. On the other hand, if the reinforcement index is 0, it is safe to discard the action and focus attention on those that remain. Eliciting additional information about the preferences regarding the connectivity of the borderline nodes can modify the number of cost-efficient portfolios and, therefore, the core index of the actions.
The core indices of the reinforcing actions (i.e., reinforcing of the individual switches) for different information sets with and without reliability requirements are depicted in Figure 5 for different budgets. There are no actions with index 1 for a small number of reinforced switches, but as they increase, the core index of some actions becomes 1. This is because a set of actions contributes similarly to the different objectives, and when just a few switches can be reinforced, it is impossible to implement all of them.
When incorporating the minimum performance requirements, the number of cost-efficient portfolios is reduced significantly, and the core indices tend to be 1 or 0. Notably, some actions that were recommendable before are discarded now, such as the reinforcement of switch 3 when the budget is five. On the other hand, some actions become relevant, such as reinforcing switches 16 and 17, even for small budgets. This highlights the importance of including the performance requirements in determining cost-efficient portfolios instead of imposing them afterward.
Figure 5: Core index for the individual reinforcement actions given different number of reinforced switches and information sets.
Centrality metrics such as degree or betweenness are usually employed to quantify the relevance of nodes in a network. For example, Ghorbani-Renani et al. [48] generate the sets of candidate nodes of interdependent water, gas, and power networks subject to protection based on five centrality metrics. One of the problems of making decisions based solely on centrality metrics is that they are not necessarily aligned with the performance of the networks. As shown before, the relevance of different switches varies with the preferences of the DM regarding the connectivity between borderline nodes or the inclusion of minimum performance requirements. Moreover, switches such as switch 10 with a good ranking according to some centrality metrics (closeness 2nd, degree 3rd, and betweenness 6th) are only relevant for specific cases and when many switches can be reinforced.
@@ 4.3.5 Impact of Estimated Failure Probability
Although failure probabilities can be estimated from historical data or reliability models, these estimates may suffer from sparse data or uncertainty. Still, it is instructive to provide recommendations based on gross estimations. We study the impact of failure probability on the composition of cost-efficient portfolios by computing them for different values of $p_{k}$ and $p^{\prime}_{k}$. The cost-efficient portfolios are the same when considering failure probabilities [^1] We do not present higher failure probabilities because the computation of the utility function is based on a cut set method, whose precision decreases for high values. Nevertheless, if required by the data, the accuracy can be improved by adding more detail to this calculation. $p_{k}\in[0.01,0.05]$ and failure probabilities after reinforcement $p^{\prime}_{k}\in\{p_{k}/2,p_{k}/3,p_{k}/4,p_{k}/5\}$.
A special case occurs if the reinforcements are perfect ($p^{\prime}_{k}=0$). The number of cost-efficient portfolios for a given $p_{k}$ decreases compared to the cases when $p^{\prime}_{k}>0$. Nevertheless, if $p^{\prime}_{k}=0$, the cost-efficient portfolios are the same for any given $p_{k}\in[0.01,0.05]$.
To understand this, consider the transportation network in Figure 6, where switches 2 and 3 have a failure probability $p_{2}=p_{3}=0.1$. The action portfolio is $q=(q_{2},q_{3})$. For the imperfect reinforcement ($p^{\prime}_{k}>0$), we obtain four cost-efficient portfolios: $(0,0),\,(1,0),\,(0,1),\,(1,1)$. However, if the reinforcement is perfect, reinforcing switch 2 or switch 3 is sufficient to guarantee that there will always be a path between the borderline nodes. Then [^2] We omit the weight in the argument of the utility function because there is a single objective., $\mathbb{E}\left[u(x)\mid(1,0)\right]=\mathbb{E}\left[u(x)\mid(0,1)\right]=\mathbb{E}\left[u(x)\mid(1,1)\right]=1$. As (1,1) is more expensive than (1,0) and (0,1), it is not cost-efficient.
Figure 6: Transportation network with borderline nodes: 1 and 4; and non-borderline nodes: 2 and 3.
This result has strong implications because it allows us to recommend portfolios of reinforcement actions even before estimating/eliciting precise values of failure probabilities. Nevertheless, accurate failure probabilities are still required to compute cost-efficient portfolios to achieve minimum performance requirements and assess the network’s performance.
@@ 5 Discussion
Our methodology assumes that individual node disruptions do not depend on the disruption of other nodes. Still, in some situations, there would be interest in modeling dependencies so that the disruption probability of one node does depend on what other nodes have been disrupted: this would be the case, for example, if the failure of a node increases the load on some other node to the extent that the disruption probability of the latter node increases. Analyses of such interdependencies could be captured through Bayesian analyses [see, e.g., 49, 50]. In some cases, incomplete information about the probability of disruption could even be admitted. For example, the disruption probabilities of the nodes could be elicited by asking experts to express statements on verbal scales and by mapping each such statement to an interval of probabilities [see, e.g., 51].
@@
Using binary variables in modeling disruptions means that the nodes are either fully operational or disrupted. To model different levels of node performance more comprehensively, one could introduce multi-valued variables to capture different gradations of node performance and, for example, to model the capacity that nodes in a transportation network have during rush hours or in the presence of minor road accidents [see, e.g., 52, 53]. A potential challenge with introducing multi-valued variables is that the number of network states would quickly grow (e.g., if three states are permitted for each node in a network containing $n$ nodes, the total number of network states would be $3^{n}$ as opposed to $2^{n}$ in a network with binary variables).
One could also extend this approach by building scenarios to characterize different external conditions in which the network must operate. For example, if the node disruption probabilities depend on the weather, specifying scenarios representing different weather conditions could be beneficial. Then, one could examine reinforcement actions to determine if a given portfolio will offer a sufficiently high level of network performance across all such scenarios or, assuming that probabilities can be associated with the different scenarios, to determine what the expected level of performance of the network is when taking the expectations over all such scenarios.
The set of cost-efficient portfolios remains invariant to changes in the component failure probabilities, given that the components have initial failure probabilities and the reinforcement actions are equal. This is relevant because uncertainty can arise due to inherent variability in these parameters, limited data, modeling inaccuracies, and unpredictable external influences.
The usability of the approach is showcased by analyzing a single station, but it can be applied to an entire railway network or many other networked systems. The railway network can be modeled by representing railway stations as nodes and railway tracks as edges connecting stations. This setting can be extended to combine the approaches, allocating resources among the railway stations and to different switches at each station.
Our illustrative example considers a single reinforcement action per node. Still, the same methodology can be used to analyze actions that affect multiple nodes simultaneously or multiple actions affecting a single node. Moreover, we focus the study on increasing the reliability of the station, but one could analyze other aspects, such as resilience, by modifying the utility functions.
@@ 6 Conclusion
In this paper, we have developed an approach for assessing the risks of infrastructure networks represented by nodes and edges that may be disrupted due to events such as natural hazards, technical failures, or intentional attacks. This assessment helps quantify the possible impact of the disruptions and identify which components (represented by nodes) should be reinforced, given multiple objectives and resource constraints. Our approach accounts for the preferences of the DM regarding the objectives and can accommodate minimum performance requirements to guarantee an adequate performance of single objectives.
Our approach opens avenues for further methodological and applied research in planning and implementing reinforcement actions for critical infrastructure systems. For example, although we have focused on a single network, the simultaneous consideration of multiple interdependent networks consisting of energy, transportation, and communication systems, especially where different decision-makers control each network separately, calls for further methodological extensions. Similarly, such an infrastructure system could handle multiple commodities (e.g., multi-commodity rail networks [54, 55]), enabling different commodity-specific DMs and objectives. Furthermore, the approach could be extended to situations where the disruption probabilities are contingent not only on the selected reinforcement actions but also on disruptions in other parts of the network or specific environmental conditions depicted by scenarios.

### v2

@@ 4.2 Preferences Regarding Connections
We consider two situations representing different specifications of preferences regarding the connections. In the first, the DM does not state preferences about the three connections. This situation is represented by the set $\mathcal{S}^{0}=\big{\{}w\in\mathbb{R}^{3}\mid w_{j}\geq 0\ \forall j,\sum_{j=1}^{3}w_{j}=1\big{\}}$.
In the second, the DM provides a ranking of the connections based on their yearly traffic volume: (A,B) at 500 trains/year, (A,C) at 500 trains/year, and (B,C) at 100 trains/year [^1] These values are provided for illustrative purposes and are not actual traffic volumes.. This is represented by the set $\mathcal{S}^{1}=\big{\{}w\in\mathbb{R}^{3}\mid\;w_{j}\geq 0\ \forall j,\;\sum_{j=1}^{3}w_{j}=1,w_{2}\geq 5w_{1},\;w_{3}\geq 5w_{1}\big{\}}$. For this last situation, we explore how the DM can ensure that the reliability of the individual connections will exceed 95% and 96%. We denote these reliability requirements by $\mathcal{S}^{1}\wedge\alpha=0.95$ and $\mathcal{S}^{1}\wedge\alpha=0.96$ respectively.
@@ 4.3.1 Cost-efficient Portfolios
We evaluated portfolios of up to 20 fortification actions, assuming that the cost of each action is the same. The corresponding cost-efficient portfolios, computed with Algorithm 1 for different information sets and minimum reliability requirements, are in Table 1. As a reference, the number of feasible portfolios is provided for different numbers of fortified switches.
Table 1: Number of cost-efficient portfolios for different information sets and reliability requirements.
Fortified switches 1 5 10 15 20
@@
When preferences about traffic volume and minimum reliability requirements are provided, te number of cost-efficient portfolios becomes much smaller. Thus, there are fewer viable portfolios of fortification actions, which makes it easier to provide recommendations about these actions.
@@ 4.3.2 Reliability of the Connections
The reliability of the connections for the cost-efficient portfolios based on the set $S^{0}$ is in Figure 3. Each point corresponds to a different portfolio for a given connection. The marginal reliability improvement gained by one additional switch decreases with the number of actions. This is because the most impactful actions are implemented first and because the actions are of equal cost. However, in general, if the actions are not of equal cost, the marginal improvement may not be decreasing because there could be very impactful actions of high cost that can be implemented only if there are sufficient resources to do so.
@@ 4.3.3 Minimum Performance for a Single Connection
The reliability of the connections for all cost-efficient portfolios containing up to five fortified switches is in Figure 4 for the cases of non-informative and incomplete information. In the case of incomplete information, there is relatively little improvement in the reliability of the connection (B, C) because this connection has less traffic than those that include the border node A.
The non-informative set contains the weights (1,0,0), (0,1,0), and (0,0,1). These weights prioritize a single connection; for example, for the weight (1,0,0), the reliability of the first connection is the only one that contributes to the utility function. Therefore, the maximum reliability that can be achieved for every single connection is given by cost-efficient portfolios for the non-informative set (see Lemma 1). For example, the maximum reliability of the connection (B, C) achieved by fortifying up to five switches is around $0.96$.
Moreover, if the minimum reliability requirement is set at $0.95$, there are some cost-efficient portfolios in the incomplete information set $S^{1}$ that do not meet this requirement, treated as a separate constraint on top of the aggregate cost of fortification actions, which is included in the characterization of otherwise feasible portfolios. If this requirement is tightened to 0.96, it can be met only by portfolios that are not cost-efficient in $S^{1}$.
The analysis of cost-efficient portfolios (two last rows of Table 1), computed with Algorithm 2 subject to the reliability requirements at 0.95 and 0.96, provides information about how many switches must be fortified to fulfill reliability requirements. For example, at least three switches must be fortified to achieve a reliability level of 0.95 and five for 0.96.
Figure 4: Reliability of connections for cost-efficient portfolios for up to five fortified switches. Values are presented for an incomplete information set ($S^{1}$) and the non-informative set ($S^{0}$).
4.3.4 Selecting Switches to Fortify
The composition of cost-efficient portfolios can be studied to derive recommendations for the selection of fortification actions. If there is a single cost-efficient portfolio at a budget level (i.e., at a given number of fortified switches), the actions in this portfolio should be selected at this budget level. However, in general, there are many cost-efficient portfolios after eliciting preference information and introducing reliability requirements. In such cases, the core index can be computed to derive recommendations [49]. This index is defined as the relative share of the action in all non-dominated portfolios. By definition, all cost-efficient portfolios with the same cost are non-dominated. We denote by $\mathcal{Q}_{ND}(c)$ the set of non-dominated portfolios of cost $c$. Then, the core index of an action $q_{k}$ when there are $c$ resources is given by
$CI(q_{k},c)=\dfrac{\lvert\{q\in\mathcal{Q}_{ND}(c)\mid q_{k}\in q\}\rvert}{\lvert\mathcal{Q}_{ND}(c)\rvert},$ (6)
@@ 4.3.4 Selecting Switches to Fortify
At a given budget level, a fortification action whose core index is 1 can be recommended because it is in every non-dominated portfolio. If the core index is 0, the action can be discarded, and the focus can be placed on the other actions. Eliciting additional preferences about the importance of connections between border nodes tends to reduce the number of cost-efficient portfolios and, therefore, change the core index of the actions, too.
Figure 5 shows the core indices of the fortification actions for switches for the two different information sets with and without reliability requirements at different budget levels. If only a few switches can be fortified, there are no actions with index 1 because these actions do not contribute to improving the reliability of all connections. At higher budget levels, which allow more switches to be fortified, some actions are in the portfolios that improve the reliability of all connections, and thus, their core index becomes 1.
If there are minimum performance requirements, there are far fewer cost-efficient portfolios that satisfy these constraints, too. Thus, the core indices tend to be 1 or 0. Some of the previously recommended actions will be discarded, such as fortifying switch 3 at a budget level of five switches. In contrast, some actions become relevant, such as fortifications of switches 16 and 17, even for small budgets. Thus, performance requirements should be introduced at the outset rather than afterward.
Figure 5: Core index for the individual fortification actions given different number of fortified switches and information sets. Red: core index 0, blue: core index 1, and gray: otherwise.
Centrality metrics such as degree or betweenness are widely used to quantify the relevance of nodes in a network. For example, Ghorbani-Renani et al. [50] examine five centrality metrics to guide the selection of nodes to be protected in interdependent water, gas, and power networks. However, a concern about using centrality metrics to guide fortification actions is that these metrics do not necessarily reflect the performance of the networks. As we have shown, the relevance of different switches depends on the DM preferences for the importance of connections and the presence of minimum performance requirements. For example, switch 10 is ranked highly by some centrality metrics (closeness 2nd, degree 3rd, and betweenness 6th), but it is only relevant for fortifying the network if many switches can be fortified.
@@ 4.3.5 Impact of Estimated Failure Probability
Although failure probabilities can be estimated from historical data or reliability models, this estimation task may be difficult due to sparse or inaccurate data. This notwithstanding, it can be instructive to provide recommendations based on approximate estimates. Therefore, we assess in the study the impact of the probability of failure on the composition of cost-efficient portfolios for different values of $p_{k}$ and $p^{\prime}_{k}$. The cost-efficient portfolios are here the same when considering failure probabilities [^2] We do not present higher failure probabilities because the computation of the utility function is based on a cut set method, whose precision decreases for high values. Nevertheless, if required by the data, the accuracy can be improved by adding more detail to this calculation. $p_{k}\in\{0.01,0.02,0.03,0.04,0.05\}$ and failure probabilities after fortification $p^{\prime}_{k}\in\{p_{k}/2,p_{k}/3,p_{k}/4,p_{k}/5\}$.
A special case occurs for perfect fortification at $p^{\prime}_{k}=0$, which means that the possibility of disruption is completely eliminated at the node where the action is implemented. In this case, for a given $p_{k}$, there are fewer cost-efficient portfolios than for $p^{\prime}_{k}>0$. If $p^{\prime}_{k}=0$, the cost-efficient portfolios are the same for $p_{k}\in\{0.01,0.02,0.03,0.04,0.05\}$.
To understand this, consider the transportation network in Figure 6, where switches 2 and 3 have a failure probability $p_{2}=p_{3}=0.1$. The action portfolio is $q=(q_{2},q_{3})$. For the imperfect fortification ($p^{\prime}_{k}>0$), there are four cost-efficient portfolios: $(0,0),\,(1,0),\,(0,1),\,(1,1)$. However, if the fortification is perfect, the fortification actions at switch 2 and switch 3 are individually sufficient to guarantee that there is a path between the border nodes. Then [^3] We omit the weight in the argument of the utility function because there is an individual objective., $\mathbb{E}\left[u(x)\mid(1,0)\right]=\mathbb{E}\left[u(x)\mid(0,1)\right]=\mathbb{E}\left[u(x)\mid(1,1)\right]=1$. As (1,1) is more expensive than (1,0) and (0,1), it is not cost-efficient.
Figure 6: Transportation network with border nodes: 1 and 4; and non-border nodes: 2 and 3.
The significance of this result is that recommendations about fortification actions can be made even if the estimated failure probabilities are not fully accurate. However, accurate estimates would be needed to assess the network’s performance or to compute cost-efficient portfolios that are guaranteed to achieve minimum reliability requirements.
@@ 5 Discussion
Our methodology assumes that individual node disruptions do not depend on the disruption of other nodes. In some situations, however, the disruption probability of a node may depend on what other nodes have been disrupted: this would be the case, for example, if the failure of a node increases the load on some other node so that the disruption probability of the latter increases. Analyses of such interdependencies could be captured through Bayesian analyses [see, e.g., 51, 52]. Challenges arising from the resulting increase in the number of required parameters can be alleviated by admitting incomplete information about disruption probabilities. This information could be obtained, for example, by asking experts to express statements on verbal scales and mapping each statement to an interval of probabilities [see, e.g., 53].
@@
Using binary variables to model disruptions assumes that nodes are operational or disrupted. To model different levels of node performance more comprehensively, one can introduce multi-valued state variables to capture different gradations of node performance and, for example, to model the capacity that nodes in a transportation network have during peak hours or in the presence of minor road accidents [see, e.g., 54, 55]. A potential challenge with multi-valued variables is that the number of network states grows quickly (e.g., if there are three rather than two states at each node in a network with $n$ nodes, the number of network states grows from $2^{n}$ to $3^{n}$).
The proposed approach can be extended to build scenarios that characterize external operating conditions. For example, specifying different weather scenarios could be instructive if the node disruption probabilities depend on the weather. This allows one to explore the robustness of cost-efficient portfolios by assessing which portfolios perform relatively well across all scenarios. Moreover, if probabilities are associated with scenarios, one can gain insight into the expected performance of the network in view of probabilities about scenarios and node disruption.
Although we have showcased the usefulness of the portfolio optimization approach by analyzing a single station, it can be applied to larger systems to study railway networks more comprehensively or even multiple interconnected networks. A critical step towards such extensions can be built through hierarchical models so that the railway network can be modeled at a higher level or aggregation by treating individual railway stations as nodes and railway tracks as edges that connect these stations. The cost-efficient portfolios for the stations can then be combined to explore how resources should be allocated among the stations to contribute to the performance objectives of the network.
Our illustrative example considers a single fortification action per node. However, the same methodology can be used to analyze choices among multiple actions at a single node or even cross-cutting actions that affect the reliability of multiple nodes. From the modeling perspective, these extensions are relatively straightforward in that they can be accommodated through logical constraints. For example, if a given cross-cutting fortification action improves the reliability of multiple nodes, then this is equivalent to our basic formulation subject to the constraint that the corresponding improvement in reliability is attained at all the nodes affected by this action. An additional extension would be to explicitly consider the sequential dynamics of implementing fortification actions. That is, if there are constraints on how many fortification actions can be implemented per unit of time within the planning horizon, the question becomes in which order the actions should be implemented so that the performance of the network can be improved as quickly as possible. One could also explore the restoring order of the nodes if the network’s performance has been compromised and the aim is to improve its resilience by restoring performance as quickly as possible.
@@ 6 Conclusion
In this paper, we have developed a multi-objective portfolio optimization approach to support the fortification of infrastructure networks whose nodes may be disrupted due to events such as natural hazards, technical failures, or intentional attacks. This optimization is based on probabilistic risk assessment in the quantification of disruption impacts and the identification of which nodes of the network should be fortified, given multiple objectives, resource constraints, and possible reliability requirements. Our approach explicitly accounts for the DM’s preferences for the different performance objectives of the network. It also accommodates minimum performance requirements regarding these objectives.
The proposed approach opens avenues for further work on strengthening critical infrastructure systems through fortification actions. A relatively straightforward extension is to consider infrastructure networks for the transportation of multiple commodities (e.g., multicommodity rail networks [56, 57]), as such commodities could be prioritized in the same way as the connections between border nodes in this paper. A somewhat more complex extension would be to consider multiple interdependent networks for energy, transportation, and communication because the dependencies between these networks would need to be modeled. The approach could be extended to situations where the disruption probabilities are contingent not only on the selected fortification actions but also on the state of other nodes or even external conditions depicted by scenarios.
Acknowledgement
This research has been partly supported by a fellow award granted by the Fulbright Finland Foundation and by the Finnish Transport Infrastructure Agency (Väylävirasto).